# What Is a Good Caption? A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

### v2

@@ 4.2 Main Evaluation Results
Table 5: The hit rate of all models. The hit rate only reflects the referring ratio of each dimension without considering the accuracy.
@@ Methods
Table 6: The $\uparrow$ (higher is better) of CAPability-QA and the result of know but cannot $\downarrow$ (lower is better) metric.
Methods
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj.
Num.
Act.
Cam.
Mov.
Event Avg.
LLaVA-OV-72B 95.0/20.3 54.6/64.3 63.8/39.6 94.0/49.6 96.2/13.9 60.6/33.0 66.3/13.7 82.0/9.4 32.1/74.8 52.2/66.1 75.5/53.2 15.7/31.4 85.3/67.1 67.2/41.3
Qwen2VL-72B 94.7/16.7 56.1/62.1 68.6/37.0 90.7/46.0 94.0/12.6 65.0/35.4 82.4/10.2 86.6/10.4 31.3/84.7 48.9/78.3 73.0/47.6 34.1/53.1 72.7/60.4 69.1/42.6
InternVL2.5-78B 95.5/19.1 56.9/57.8 67.1/35.4 90.0/45.2 91.2/11.4 54.1/21.4 79.5/11.0 82.5/47.0 19.1/8.2 49.7/73.0 79.1/62.4 23.3/68.1 81.7/69.9 66.9/40.8
Qwen2.5VL-72B 92.7/15.3 58.2/60.7 67.4/33.5 84.4/37.2 88.7/9.1 63.9/24.3 87.4/5.9 87.3/8.5 33.4/47.4 41.4/66.2 75.8/49.3 39.5/38.2 85.8/61.1 69.7/35.1
GPT-4o (0806) 94.5/13.1 47.2/55.2 72.5/26.6 79.5/34.7 84.5/7.8 71.6/16.1 80.5/6.7 79.3/3.5 37.2/30.9 46.2/64.8 81.1/41.7 20.5/53.9 78.6/51.7 67.2/31.3
Gemini-1.5-pro 97.3/11.9 51.6/41.0 78.8/24.1 94.4/36.4 87.1/9.6 56.8/19.2 84.8/5.5 84.2/3.1 41.2/18.0 51.2/51.6 74.4/36.1 32.2/23.4 82.8/47.9 70.5/25.2
Gemini-2.0-flash 98.3/16.3 46.8/52.6 73.3/32.8 93.4/45.1 95.2/12.6 57.6/9.0 84.8/4.5 74.5/3.2 49.1/25.7 44.2/58.4 81.6/45.9 24.8/23.1 86.6/54.5 70.0/29.5
F1-score of 7B models. The precision, recall, and F1-score of 7B models are shown in Tab. 4. Among these 6 models, Qwen2.5VL-7B achieves the best precision (75.5%), recall (48.1%), and F1-score (54.8%), demonstrating its awesome ability. Averagely, the 7B models perform a bit worse than 72B models, verifying the scaling law. Among the dimensions, they follow a similar pattern as 72B and closed-source models. Researchers should focus on the thoroughness of object number, object color, spatial relation, style, character identification, and event, and try to improve the ability of the action and the two camera-related dimensions.
Hit rate among all models. We also report the hit rate of these models, which represents the pure thoroughness of each dimension, as shown in Tab. 5. For example, it is considered a hit if the caption mentions any object for the object category dimension, or mentions any angle information for camera angle dimension, but for the object number or color dimension, it is only considered a hit if the caption mentions any number or color information of the correct object. We find the hit rate seems to increase as the size of models increases, which may be due to more knowledge and stronger instruction following ability for larger models. Among all dimensions, the hit rate of character identification performs the worst, the existing models prefer to keep silent as they usually cannot recognize the person and character well. The closed-source models would be more likely to tell the name of characters, and we guess this may be due to stronger knowledge and more diverse training data.
QA-based evaluation and the $\bm{K\bar{T}}$ metric. As we convert our annotations to QA format, we evaluate the accuracy of closed-source APIs and 72B models, thus further calculating their $K\bar{T}$ metrics, as shown in Tab. 6. We are surprised to find that the difference in their QA accuracy is not significant, which means they can have a similar level of understanding of the correct visual descriptions based on the specific questions. When it comes to the $K\bar{T}$ metric, these models vary a lot. Gemini-1.5-pro performs the best with the smallest $K\bar{T}$ (25.7%), but the $K\bar{T}$ of LLaVA-OneVision-72B, Qwen2VL-72B, and InternVL2.5-78B comes to more than 40%, which means they are more likely knowing the answer but cannot tell automatically. This phenomenon shows the performance gap between the strong instruction (QA) task and the weak instruction (caption) task, which may be ignored by previous work.
4.3 Stability Analysis
To validate the stability and robustness of our GPT-4-Turbo based evaluation method, we take the inferred caption of Gemini-1.5-pro as the example, run our evaluation 5 times, and the result is shown in Fig. 7. We tag the fluctuation range, i.e., the difference between the maximum and minimum scores, besides the data point. Fig. 7 shows our strong stability, and our average range of precision and recall are 1.1% and 1.0%, respectively. This demonstrates the reliability and interpretability of our evaluation method, which matches annotated elements in the generated captions. See Sec. A.2 for more stability analysis.
5 Conclusion
In this work, we present CAPability, the first comprehensive benchmark for evaluating visual caption across images and videos through 6 views and 12 dimensions analysis. Unlike existing benchmarks that rely on oversimplified metrics or limited visual elements, CAPability introduces a correctness-thoroughness dual evaluation framework based on precision, recall, hit rate, and $K\bar{T}$. Through this meticulous evaluation process, we uncover specific strengths and areas needing improvement across leading models, such as their proficiency in object counting and challenges in aspects like camera angle detection, character identification and action recognition. We also indicate the ”know but cannot tell” phenomenon of MLLMs, which may be ignored by previous work. We believe that CAPability will play a pivotal role in advancing the field of visual captioning by encouraging the development of models that holistically understand and describe visual content. We open-source all our annotated data to facilitate future research.
Appendix
Appendix A More Experimental Analysis
A.1 Implementation Details
For all our evaluated model, we follow their official configurations to run the inference. We set the temperature of all open-source models to 0, while keeping the default for closed-source APIs. All maximum output token length is set to 8192. We list the configurations as follows.
NVILA. We use the official image and video process function in VILA repository, and uniformly sample 8 frames for videos, as it is suggested in the official config.
Gemini-1.5-pro and Gemini-2.0-flash. As Gemini API supports video, we directly send the original image and video to the API server. For very few videos with too large file size, we downsample the fps into 3, and send the downsampled video to the API server for connection stability.
A.2 More Stability Analysis
@@
To further evaluate the stability of our evaluation pipeline, we conduct another experiment. Specifically, we introduce two more evaluation models, Gemini-2.0-flash and Qwen2.5-Max, as they are both the most popular and powerful SOTA language models. We re-run the evaluation process with new evaluation models with the same evaluation prompts. The average result of evaluations for closed-source models and open-source 72B models is shown in Fig. A1. Though the judgment criteria for each model may be slightly different, leading to only a slight difference in the scores, but the evaluation results show high consistency as they keep the same rank of these MLLMs as GPT-4-Turbo. This demonstrates the high reliability, interpretability, and stability of our evaluation methods.
Appendix B More Benchmark Details
B.1 Details of Dimensions
We explain each dimension in detail about what they represent here.
Object category. This dimension measures the ability of whether models can give a correct description about a specific object in the image. The object is randomly selected from the image.
Object color. Given a kind of randomly selected object in an image, this dimension measures the ability of whether models can describe the color correctly.
Spatial relation. Given two nearby objects in an image, this dimension measures the ability of whether models can describe the spatial relationship of the two objects correctly. We sample 500 images from our collected data, and sample another 500 images from CompreCap, with their spatial relationship descriptions.
Action. Given a video, this dimension measures the ability of whether models can recognize the action in the video. We use the video data of Dream-1K and re-annotate the action from their annotations.
Event. Given a video, this dimension measures the ability of whether models can tell a complete event in the video. We refer Dream-1K to design this dimension, and we extract the events from their annotations. Different from other dimensions with atom-level elements, the event is usually composed of subjects and actions, which measures the temporal summarization ability of the model.
B.2 Prompts of Inference and Evaluation
Evaluation prompt. As we divide the evaluation of dimensions into two types: 1) dimensions with specific categories (i.e., style, camera angle, and camera movement), 2) dimensions with open-ended descriptions. Therefore, we design two kinds of templates for evaluating, and fine-tune them within each dimension. In Fig. A3, we take the object number dimension and camera movement dimension as examples, to show our prompts for evaluation. For dimensions with specific categories, we ask GPT-4-Turbo to extract the key information and classify the caption into our pre-defined categories or the ’N/A’ class. The correct classification is considered as positive, the wrong one as negative, and the ’N/A’ result is considered as a miss. For dimensions with open-ended descriptions, we ask GPT-4-Turbo to directly compare the annotations and the caption, and give out the result of positive, negative, or miss with reasons.
B.3 Explanation for One Represents All Strategy
For each dimension, we only annotate one element, though there may be more than one element existing for some dimensions. Therefore, our annotations do not cover the whole visual content. But for those who try to cover the whole visual content, it is actually pretty hard to achieve this, as we mentioned in Sec. 3.1, the objects can be divided into almost infinite granularity. We focus on keeping the randomness of elements selection, thus covering the whole visual content in a statistical sense, based on the law of large numbers. Therefore, we get the ability to evaluate the thoroughness of the generated captions by calculating the recall and hit rate.
B.4 Benchmark Examples
Examples of converted QA pairs. As we directly annotate the visual elements in the image or video rather than the caption sentence, we can easily convert our annotation into the format of question-answer (QA) pairs, and we name it as CAPability-QA. We use CAPability-QA to evaluate the QA accuracy and the know but cannot tell ($K\bar{T}$) metric. In Fig. A5, we also show the same visual cases with Fig. A4 for each dimension with converted QA format. Most of the dimensions are converted to the format of a multiple-choice QA task with several options, and the object color, OCR, and character identification dimensions are designed as open-ended QA tasks.
Examples of inference and evaluation. In Fig. A6 and Fig. A7, we visualize the inferred caption of Gemini-1.5-pro, GPT-4o (0806), and Qwen2.5VL-72B in object number dimension and camera angle dimension. In Fig. A6, the annotation of the given image is 7 pig trotters. Gemini-1.5-pro refers to the correct number of pig trotters, and we thus give it a positive. GPT-4o recognizes the trotters in the image, but counts with a wrong number, 6, and we thus give it a negative. As for Qwen2.5VL-72B, it says there are 7 chicken thighs in the image, recognizing the wrong object category. However, this is the dimension of the object number, and therefore we only evaluate the correctness and thoroughness of the number, without considering the categories of objects. As Qwen2.5VL-72B does not mention the pig trotters, we give it a miss. In Fig. A7, the annotation of the given image is dutch angle, which means the lens has a certain angle of deflection along the central axis, making the horizon crooked. Gemini-1.5-pro says the image is taken from a medium-high angle, and we classify it into the high angle category, thus negative. GPT-4o explicitly points it out as a subtle dutch angle, thus is classified into the dutch angle category, which is positive. Qwen2.5VL-72B describes the image shot from a slightly elevated angle, and it appears to be a level angle, which is also negative. These two figures show our evaluation pipeline, which is precise and reliable.

### v3

@@ B.3 More Experimental Results
Table A1: The referring ratio of all models, which only reflects the referring ratio of each dimension without considering the accuracy.
@@ Methods
Figure A7: The evaluation of repeating 5 times for Gemini-1.5-pro captions. We tag the fluctuation range beside the data point.
@@
Referring ratio among all models. Apart from the precision and hit, we can also report another metric, referring ratio, which represents the referring ratio about the dimension in visual caption and can be calculated as:
Referring $\text{Referring Ratio}=\frac{|S(\text{COR})\cup S(\text{INC})|}{|S(\text{ALL})|}.$ (A5)
This metric only considers the pure thoroughness of the caption in each dimension without considering the accuracy.
We report the referring ratio in Tab. A1. For example, it is considered a reference if the caption mentions any object for the object category dimension, or mentions any angle information for the camera angle dimension, but for the object number or color dimension, it is only considered a reference if the caption mentions any number or color information of the correct object. We find the referring ratio seems to increase as the size of models increases, which may be due to more knowledge and stronger instruction following ability for larger models. Among all dimensions, the referring ratio of character identification performs the worst, the existing models prefer to keep silent as they usually cannot recognize the person and character well. The closed-source models would be more likely to reveal the names of characters, and we guess this may be due to stronger knowledge and more diverse training data.
Evaluation stability. To validate the stability and robustness of our GPT-4-Turbo based evaluation method, we take the inferred caption of Gemini-1.5-pro as the example, run our evaluation 5 times, and the result is shown in Fig. A7. We tag the fluctuation range, i.e., the difference between the maximum and minimum scores, besides the data point. Fig. A7 shows our strong stability, and our average range of precision and recall are 1.1% and 1.0%, respectively. This demonstrates the reliability and interpretability of our evaluation method, which matches annotated elements in the generated captions.
@@
Moreover, we introduce two other models, Gemini-2.0-flash and Qwen2.5-Max, to replace the GPT-4-Turbo in our evaluation pipeline. Both of them the most popular and powerful SOTA language models. We re-run the evaluation process with new evaluation models with the same evaluation prompts. The average result of evaluations for closed-source models and open-source 72B models is shown in Fig. A8. Though the judgment criteria for each model may be slightly different, leading to only a slight difference in the scores, but the evaluation results show high consistency as they keep the same rank of these MLLMs as GPT-4-Turbo. This further demonstrates the high reliability, interpretability, and stability of our evaluation methods.
@@
B.4 Visualization of Inference and Evaluation.
In Fig. A9 and Fig. A10, we visualize the inferred caption of Gemini-1.5-pro, GPT-4o (0806), and Qwen2.5VL-72B in object number dimension and camera angle dimension. In Fig. A9, the annotation of the given image is 7 pig trotters. Gemini-1.5-pro refers to the correct number of pig trotters, and we thus give it a positive. GPT-4o recognizes the trotters in the image, but counts with a wrong number, 6, and we thus give it a negative. As for Qwen2.5VL-72B, it says there are 7 chicken thighs in the image, recognizing the wrong object category. However, this is the dimension of the object number, and therefore we only evaluate the correctness and thoroughness of the number, without considering the categories of objects. As Qwen2.5VL-72B does not mention the pig trotters, we give it a miss. In Fig. A10, the annotation of the given image is dutch angle, which means the lens has a certain angle of deflection along the central axis, making the horizon crooked. Gemini-1.5-pro says the image is taken from a medium-high angle, and we classify it into the high angle category, thus negative. GPT-4o explicitly points it out as a subtle dutch angle, thus is classified into the dutch angle category, which is positive. Qwen2.5VL-72B describes the image shot from a slightly elevated angle, and it appears to be a level angle, which is also negative. These two figures show our evaluation pipeline, which is precise and reliable.
Appendix C Copyright and License
CAPability comprises data from SA-1B, COYO-700M, Wukong-100M, Wikipaintings, VSI-Bench, CompreCap, Dream-1K, and uses some craweled data from inner retrieval system, each governed by its own licensing terms:
  • SA-1B: SA-1B Dataset Research License [^1] https://ai.meta.com/datasets/segment-anything-downloads/
  • COYO-100M: CC-BY-4.0 License
  • Wukong-100M: CC-BY-NC-SA-4.0 License
  • Wikipaintings: BSD 2-clause License
  • VSI-Bench: Apache-2.0 License
  • CompreCap: CC-BY-4.0 License
  • Dream-1K: Apache-2.0 License
  • Craweled data: Our craweled data are all retrieved from inner multi-modal retrieval system, which contains various public datasets, and visual contents from websites with CC0 license.
Appendix D Limitations
Different from multi-choice VQA benchmarks which evaluate models by definite and explicit choice accuracy, our CAPability is a visual caption benchmark, which still depends on LLMs for evaluation. Therefore, it is still hard to ensure a completely correct evaluation. We try to split the evaluation into several dimensions, thus makes the evaluation as simple and clear as possible. Therefore, the LLM-based evaluation can be more accurate. Due to the LLM language capability limitation, it can still make wrong judgments and requires a carefully designed prompt for constraint.
Appendix E Societal Impacts
As our proposed CAPability can perform comprehensive caption evaluations of MLLMs, this work can help LLM users make informed choice, and leads the community to build more and more strong MLLMs. The potential negative impacts are similar to other LLM-related works, The development of MLLMs and MLLMs’ benchmarks pose societal risks like the perpetuation of biases, the potential for misinformation, job displacement, etc.